OpenStudy (anonymous):

Find u*v u={-2,-7,-2} v={-3,6,-5}. VECTORS

OpenStudy (anonymous):

There are two ways to multiply 3D vectors. \[ \mathbf u\cdot \mathbf v \\ \mathbf u \times \mathbf v \]

OpenStudy (anonymous):

With the x

OpenStudy (anonymous):

You should learn how to do the cross product.

OpenStudy (anonymous):

Enlighten me if you don't mind on the cross product

OpenStudy (anonymous):

It's basically a determinant. Given two vectors \(\vec{u}=\langle u_1,u_2,u_3\rangle=u_1\vec{i}+u_2\vec{j}+u_3\vec{k}\) and \(\vec{v}=\langle v_1,v_2,v_3\rangle\), the cross product of \(\vec{u}\) and \(\vec{v}\) is \[\vec{u}\times\vec{v}=\begin{vmatrix} \vec{i}&\vec{j}&\vec{k}\\ u_1&u_2&u_3\\v_1&v_2&v_3\end{vmatrix}=(u_2v_3-u_3v_2)\vec{i}-(u_3v_1-u_1v_3)\vec{j}+(u_2v_2-u_2v_1)\vec{k}\] (\(\vec{i},\vec{j},\vec{k}\) are the standard unit vectors) There are plenty of online resources you can refer to if you don't quite understand it right away.

OpenStudy (zzr0ck3r):

generally \[u*v=u\cdot v\]